Break-Even Analysis Pinnacle Party Inflatables, a supplier of inflatable bouncy houses, has budgeted the following amounts for its next fiscal year:

Total fixed expenses $238,500

Selling price per unit $290

Variable expenses per unit $140

If Pinnacle Party Inflatables can reduce fixed expenses by $47,250, by how much can variable expenses per unit increase and still allow the company to maintain the original break-even sales in units? Round answers to two decimal places, if applicable.
Answer #1

Solution

Information given:

Total Fixed Expenses = $ 238500

Selling Price per unit = $ 290

Variable Expenses per unit = $ 140

Contribution per unit = Selling Price per unit - Variable Expenses per unit

= $ (290 - 140) = $ 150

Therefore, currently Break-Even Point (Units) will be = Total Fixed Costs / Contribution per unit

= 238500 / 150

= 1590 units

Now, the company wants to reduce its fixed costs by $ 47,250, or $ (238500 - 47250) = $ 191250 by maintaining the same Break-Even point,

Therefore, Fixed Costs / Contribution per unit = 1590

Or, 191250 / Contribution per unit = 1590

Or, Contribution per unit = 191250 / 1590

Or, Contribution per unit = 120.28 (Approx)

Again,

Selling Price per unit - Variable Cost per unit = Contribution per unit

Or, 290 - Variable Cost per unit = 120.28

Or, Variable Cost per unit = 290 - 120.28

Or, Variable Cost per unit = 169.72

Therefore,

Existing Variable Cost per unit = $ 140

Revised Variable Cost per unit = $ 169.72

Increase in Cost per unit = $ (169.72 - 140) = $ 29.72

Answer: Variable Cost per unit can be increased by $ 29.72 allowing the company to maintain same break-even point after reducing fixed costs by $ 47,250.